# AutothinkRAG: Complexity-Aware Control of Retrieval-Augmented Reasoning for Image-Text Interaction

### 05 Mar 2026 v1

@@ Abstract
Information-intensive Document Question Answering (DocQA) is often constrained by long contexts and information overload, which hinders Vision-Language Models (VLMs) from performing precise direct reasoning. Although multimodal GraphRAG has achieved preliminary breakthroughs, existing frameworks still face dual challenges: (1) the necessity of large-scale models for handling queries of diverse complexities and (2) the inherent reasoning bottlenecks of end-to-end VLMs. To address these issues, we propose AutoThinkRAG, a framework that enhances the understanding of complex documents by synergizing the capabilities of multiple models. Specifically, we introduce a Query Complexity Router to allocate reasoning paths based on the analysis of query difficulty. Furthermore, to overcome the reasoning boundaries of VLM, we propose a functional decoupling architecture: a small-scale VLM serves as a high-fidelity visual interpreter to transform query-relevant visual cues into textual representations, which are subsequently processed by an LLM for logical deduction and synthesis. Extensive experiments on DocBench and MMLongBench demonstrate that AutoThinkRAG significantly reduces inference costs while achieving new state-of-the-art performance. Further ablation studies verifies the effectiveness of our proposed method.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) serves as a core technology for processing heterogeneous multimodal documents, including financial PDFs and technical diagrams. Its objective is to retrieve multimodal information relevant to a given query from documents as contextual content for question answering, thereby enhancing the accuracy and richness of generated results. As an evolving direction in this field, Multimodal RAG (MM-RAG) needs to address two core issues simultaneously: multimodal information processing and high-precision reasoning.
@@
Knowledge graphs can effectively represent the relationships and attributes of information, while leveraging dense vector representations to identify semantically relevant content lacking direct structural connections. In recent years, researchers have conducted extensive explorations on multimodal graph RAG and complex question answering. As the state-of-the-art (SOTA) approach, RAGAnything (Guo et al., 2025) has validated the application of multimodal graphs in document question answering. HyperGraphRAG (Luo et al., 2025) effectively captures high-order dependencies in complex knowledge bases through hyperedge modeling, improving the accuracy of complex multi-hop reasoning. MMGraphRAG (Wan and Yu, 2025) provides a new technical paradigm for multimodal RAG in graph RAG scenarios.
@@
Despite the progress achieved by existing methods, current SOTA frameworks still face two critical issues that severely impact performance:(1) Retrieval Rigidity: Existing systems adopt a static retrieval strategy that fails to make reasonable judgments and planning based on query complexity. Accurate understanding of queries with arbitrary complexity requires large-scale models, and this high-computation-cost approach leads to the inefficient allocation of computing resources.(2) Reasoning Deficit: Multimodal information question answering relies on monolithic Vision-Language Models (VLMs) for end-to-end generation. However, recent related studies have shown that VLMs exhibit significantly lower reasoning performance than Large Language Models (LLMs), which consequently results in the problem of "correct visual recognition but incorrect answer generation"(Wan and Yu, 2025).
@@
In this paper, we propose the AutoThink-RAG framework, which addresses the aforementioned issues through modular cognitive orchestration. To tackle retrieval rigidity, we design the AutoThink lightweight cognitive router, which dynamically selects retrieval paths based on query parsing and complexity evaluation, enabling on-demand allocation of computing resources. To resolve reasoning deficit, we propose the Decomposition of Perception and Reasoning (DPR) architecture, which limits the role of VLMs to that of "visual translators" and delegates logical integration tasks to LLMs, achieving more accurate reasoning without loss of relevant information.
Our contributions are three-fold:
@@
  • We present AutoThink-RAG, a scalable architecture that integrates MinerU-based parsing with a hybrid Graph-Vector storage, establishing a new pareto-optimal frontier between efficiency and accuracy;
@@ 3.1 Overall Workflow
AutoThinkRAG deconstructs the multimodal Document Question Answering (DocQA) task into intent-aware routing and decoupled reasoning. The end-to-end inference process is formally defined as:
@@ 3.2.3 Hybrid Storage and Transmission Protocol
    Vector Store: All entities, relations, and text chunks are mapped into a dense embedding space, with $S,B,$ and $P$ indexed as key attributes.
Information Transmission Mechanism. The framework achieves efficient cross-module transmission via a metadata-driven protocol. During retrieval, the storage path $S$ enables the direct transmission of raw multimodal data to the VLM. Simultaneously, leveraging $B$ and $P$, the system fetches adjacent context from the original document, effectively bridging the “Information Gap” between isolated fragments and their original document context.
Doc Types Domains All
Aca. Fin. Gov. Law. News Txt. Mm. Una.
LLM 41.3 16.3 50.7 49.7 77.3 53.9 20.1 75.8 44.8
MLLM 19.8 16.3 28.4 31.4 46.5 35.7 15.9 39.5 27.7
$M$ Locations Formats Overall
Sin. Mul. Una. C.T. Txt. Lay. Fig. All
LLM 22.5 20.0 53.2 14.7 33.3 23.5 16.1 27.8
MLLM 13.3 7.9 13.9 8.4 10.7 11.8 11.4 11.6
Figure 1: Overview of our AutoThink–RAG. The framework consists of four primary stages: (1) Information Extraction: Heterogeneous multimodal data are ingested. (2) Data Processing: Through layout analysis, the system partitions data into text chunks and image indices, followed by an iterative “Extract–Judge” loop to construct a Graph Knowledge Base. (3) Data Linker: A hybrid retrieval mechanism integrates traditional Top–K vector search with graph–based merging to capture complex entity relationships. (4) Reasoning & Response: The Query Router decomposes the original query into sub-queries ($Q_{1},Q_{2},Q_{3}$), which are then processed through a decoupled perception reasoning architecture leveraging both “related text” and visual cues to synthesize the final response.
@@ 3.3.1 Query Complexity Router (QCR)
Based on these, QCR assigns a complexity label $c(Q)$ and generates routing instructions $I_{p}$.
@@ 3.3.2 Functional Decoupling via DPR
Recent studies indicate that VLMs exhibit significantly lower capabilities in logical reasoning compared to LLMs (as shown in Table 1). This discrepancy stems from the VLM architecture’s over-reliance on intermediate projection layers for modality alignment, which struggles to maintain logical consistency in complex derivations. Conversely, VLMs demonstrate superior performance in descriptive tasks post-perception. Consequently, we propose the Decoupled Perception and Reasoning (DPR) strategy, partitioning the task into two specialized sub-processes:
@@
Method Domains Types Overall
Aca. Fin. Gov. Law. News. Txt. Mm. Una.
AutothinkRAG 80.16 75.40 85.71 87.09 92.50 86.43 77.31 81.25 82.13
Nothink 77.34 73.46 80.59 85.29 88.78 84.08 76.12 77.84 79.37
VLM only 77.03 74.00 81.54 84.39 89.26 85.32 75.09 76.42 79.47
baseline 76.34 72.93 79.77 84.07 84.52 84.65 74.47 75.83 78.02
Raganything 74.90 70.68 77.41 83.06 81.67 84.21 76.36 52.80 75.47
Method Domains Overall
Res. Tut. Acad. Guid. Broch. Admin. Fin.
AutothinkRAG 44.86 52.20 52.26 52.34 51.00 62.19 56.14 51.29
Nothink 43.15 48.55 43.21 52.25 46.00 53.08 50.42 47.21
VLM only 41.43 49.27 43.43 52.25 41.00 54.32 48.72 45.55
baseline 40.41 47.82 42.42 50.96 42.00 51.85 46.15 44.86
8B 42.80 44.12 43.21 52.34 46.00 50.61 50.42 46.09
Raganything 41.09 41.30 44.22 45.16 40.00 54.32 52.13 44.36
@@ 4 Experiments
In this section, we empirically validate the effectiveness of AutoThink-RAG. Our goal is to address two primary research questions: (1) Compared to static retrieval, can the AutoThink router effectively reduce hallucinations and handle complex query intents? (2) Does the Decoupled Perception-Reasoning (DPR) architecture demonstrate superior performance in reasoning tasks involving multimodal documents of varying lengths?
@@ 4.2 Implementation Details
In our experimental evaluation of AutoThink-RAG, we implement a multi-model collaborative framework characterized by a decoupled architecture. Specifically, we employ Qwen2.5-VL-3B as the AutoThink Router to facilitate global task orchestration and intent scheduling. For the knowledge construction phase, Qwen3-8B (non-reasoning version) is utilized to perform robust entity and relation extraction. To optimize retrieval precision, we adopt Qwen3-embedding-8B (with 1024-dimensional embeddings) for dense vector representation, followed by bge-reranker-v2-m3 for fine-grained reranking. To realize the Decoupled Perception-Reasoning (DPR) paradigm, we deploy Qwen2.5-VL-7B as the Perception Expert, which distills complex visual information into dense textual descriptions. Subsequently, Qwen3-32B-Instruct serves as the Reasoning Expert to execute high-level logical deductions based on the integrated context. Regarding hyperparameter configurations, all generative models are set with a temperature of $0.7$ and a Top-$p$ value of $0.9$. During the retrieval stage, we set the Top-$K$ for entity recall to $40$ and the number of retrieved text chunks to $20$ to ensure sufficient context coverage.
@@ 4.3 Main Results
Table 2 summarizes the performance on DocBench. AutoThink-RAG achieves a state-of-the-art Overall Accuracy of 82.13%, significantly outperforming the baseline (78.02%).
Crucially, our framework demonstrates a substantial breakthrough in the Unanswerable (Una) category, boosting accuracy from 52.80% to 81.25% (a +28.45% improvement). This result strongly validates the cognitive awareness of our AutoThink router. Unlike static retrieval systems that force-feed irrelevant context to the generator, AutoThink can detect insufficient information and guide the reasoning engine to refuse answering, thereby drastically reducing hallucinations. Additionally, in information-dense domains like News (+10.83%) and Government (+8.30%), our Hybrid Hypergraph Retrieval proves effective in capturing complex entity relationships that simple vector search misses.
@@
Table 3 presents the results on long-context tasks. AutoThink-RAG consistently surpasses the baseline across all document categories, achieving an Overall Accuracy of 51.29% (+6.43%).
The unified baseline often struggles with "contextual entrainment" in long documents, where visual noise distracts the VLM’s reasoning. In contrast, our DPR Architecture maintains high performance even in complex categories like Admin (+10.34%) and Finance (+9.99%). By converting visual signals into dense text descriptions and offloading the synthesis to a specialized LLM, our approach effectively bridges the "Reasoning Gap," allowing for robust multi-hop deduction over extensive contexts.
@@ 4.4 Baselines
To ensure experimental rigor and fairness, we built a baseline model and employed the identical model on RAGAnything (Guo et al., 2025) for testing and comparison.In the baseline configuration, we bypass query decomposition and employ the VLM for direct end-to-end inference, rather than using our proposed decoupled paradigm. For the RagAnything benchmark, we substitute its internal components with the specific models used in our framework and maintain identical hyperparameter configurations to ensure a controlled comparison.
@@ 4.5 Ablation Study
Our decoupling module transforms the direct reasoning of the VLM into a hybrid VLM+LLM indirect reasoning paradigm. To evaluate its efficacy, we conduct an ablation study by replacing this module with a standalone LLM reasoning component while keeping the upstream router constant. We evaluate the accuracy across various page ranges, as illustrated in Figure 2. The results indicate that direct VLM reasoning substantially underperforms compared to our decoupled approach, with the performance gap becoming more pronounced as the document length increases.
@@ 5 Conclusion
While Auto-thinkRAG achieves significant performance gains over existing multimodal GraphRAG architectures using models with fewer parameters, it still relies on a sequential document parsing and embedding pipeline, which limits the overall processing speed. Our future work will be dedicated to achieving the efficient coupling of document parsing and information encoding to improve the overall processing speed and accuracy.

### 17 Mar 2026 v2

@@ Abstract
Multimodal document question answering requires retrieving dispersed evidence from visually rich long documents and performing reliable reasoning over heterogeneous information. Existing multimodal RAG systems remain limited by two bottlenecks: static retrieval that ignores query complexity, and end-to-end Vision-Language Models (VLMs) that couple visual perception with logical reasoning, leading to inefficient computation and unstable answer generation. We propose AutoThinkRAG, a complexity-aware inference architecture for multimodal document QA. It has two components: (1) a Query Complexity Router that analyzes query difficulty and structure to adaptively select retrieval and reasoning paths; and (2) a Perception–Reasoning Decoupling architecture that uses a lightweight VLM as a high-fidelity visual interpreter to convert query-relevant visual cues into textual representations, which are then passed to an LLM for logical reasoning and answer synthesis. This design improves both efficiency and robustness, especially on long-document and unanswerable queries. Experiments on DocBench and MMLongBench show that AutoThinkRAG achieves 82.13% and 51.29% overall accuracy, respectively, while reducing per-query token consumption by 18.9% and monetary cost by 18.2%. Further analyses show that the gains are most pronounced on complex queries requiring adaptive retrieval and multi-step reasoning.
@@ 1 Introduction
Figure 1: Compared with traditional RAG, AutothinkRAG enables more effective and robust long-document question answering through adaptive retrieval and decoupled perception-and-reasoning.
@@
Retrieval-Augmented Generation (RAG) serves as a core technology for processing heterogeneous multimodal documents. Its objective is to retrieve multimodal information relevant to a given query from documents as contextual content for question answering, thereby enhancing the accuracy and richness of generated results. As an evolving direction in this field, Multimodal RAG (MM-RAG) needs to address two core issues simultaneously: multimodal information processing and high-precision reasoning.
@@
Knowledge graphs can effectively represent the relationships and attributes of information, while dense vector representations can identify semantically relevant content lacking direct structural connections. In recent years, researchers have conducted extensive explorations on multimodal graph RAG and complex question answering. As the state-of-the-art (SOTA) approach, RAG-Anything (Guo et al., 2025) has validated the application of multimodal graphs in document question answering. HyperGraphRAG (Luo et al., 2025) effectively captures high-order dependencies in complex knowledge bases through hyperedge modeling, improving the accuracy of complex multi-hop reasoning.
@@
Despite the progress achieved by existing methods, current SOTA frameworks still face two critical issues that severely impact performance: (1) Retrieval Rigidity: Existing systems adopt a static retrieval strategy that fails to make reasonable judgments and planning based on query complexity. Accurate understanding of queries with arbitrary complexity requires large-scale models, and this high-computation-cost approach leads to inefficient allocation of computing resources. (2) Reasoning Deficit: Multimodal information question answering relies on monolithic Vision-Language Models (VLMs) for end-to-end generation. However, recent studies have shown that VLMs exhibit significantly lower reasoning performance than Large Language Models (LLMs), leading to the problem of “correct visual recognition but incorrect answer generation” (Wan and Yu, 2025).
@@
In this paper, we propose the AutothinkRAG framework to address the aforementioned issues. To tackle retrieval rigidity, we design the Autothink lightweight cognitive router, which dynamically selects retrieval paths based on query parsing and complexity evaluation, enabling on-demand allocation of computing resources. To resolve reasoning deficit, we propose the Decomposition of Perception and Reasoning (DPR) architecture, which limits the role of VLMs to that of “visual descriptors” and delegates logical integration tasks to LLMs, achieving more accurate reasoning without loss of relevant information.
Our contributions are four-fold:
@@
  • We present AutothinkRAG, a scalable architecture that integrates MinerU-based parsing with a hybrid Graph-Vector storage, establishing a new pareto-optimal frontier between efficiency and accuracy;
@@ 2.1 Graph-based Retrieval-Augmented Generation
Figure 2: Overview of our AutothinkRAG. The framework consists of four primary stages: (1) Information Extraction: Heterogeneous multimodal data are ingested. (2) Data Processing: Through layout analysis, the system partitions data into text chunks and image indices, followed by an iterative “Extract–Judge” loop to construct a Graph Knowledge Base. (3) Data Linker: A hybrid retrieval mechanism integrates traditional Top–K vector search with graph–based merging to capture complex entity relationships. (4) Reasoning & Response: The Query Router decomposes the original query into sub-queries ($Q_{1},Q_{2},Q_{3}$), which are then processed through a decoupled perception reasoning architecture leveraging both “related text” and visual cues to synthesize the final response.
@@ 3.1 Overall Workflow
AutothinkRAG deconstructs the multimodal Document Question Answering (DocQA) task into intent-aware routing and decoupled reasoning,as illustrated in Figure  2. The end-to-end inference process is formally defined as:
@@ 3.2.3 Hybrid Storage and Transmission Protocol
    Vector Store: All entities, relations, and text chunks are mapped into a dense embedding space, with $S,B,$ and $P$ indexed as attributes.
Information Transmission Mechanism. The framework enables efficient cross-module transmission via a metadata-driven protocol. During retrieval, storage path $S$ allows direct raw multimodal data transmission to the VLM. Meanwhile, leveraging $B$ and $P$, the system fetches adjacent context from the original document, effectively bridging the “Information Gap” between isolated fragments and their source context.
@@ 3.3.1 Query Complexity Router (QCR)
Based on these, QCR assigns a complexity label $c(Q)$ and generates routing instructions $I_{p}$.Concretely, queries relying on localized evidence are regarded as Simple, whereas queries requiring multi-entity, multimodal aggregation and multi-step reasoning are regarded as Complex.
Doc Types Domains All
Aca. Fin. Gov. Law. News Txt. Mm. Una.
LLM 41.3 16.3 50.7 49.7 77.3 53.9 20.1 75.8 44.8
MLLM 19.8 16.3 28.4 31.4 46.5 35.7 15.9 39.5 27.7
$M$ Locations Formats Overall
Sin. Mul. Una. C.T. Txt. Lay. Fig. All
LLM 22.5 20.0 53.2 14.7 33.3 23.5 16.1 27.8
MLLM 13.3 7.9 13.9 8.4 10.7 11.8 11.4 11.6
Method Domains Types Overall
Aca. Fin. Gov. Law. News. Txt. Mm. Una.
AutothinkRAG 80.16 75.40 85.71 87.09 92.50 86.43 77.31 81.25 82.13
Nothink 77.34 73.46 80.59 85.29 88.78 84.08 76.12 77.84 79.37
VLM only 77.03 74.00 81.54 84.39 89.26 85.32 75.09 76.42 79.47
baseline 76.34 72.93 79.77 84.07 84.52 84.65 74.47 75.83 78.02
RAG-Anything 74.90 70.68 77.41 83.06 81.67 84.21 76.36 52.80 75.47
Method Domains Overall
Res. Tut. Acad. Guid. Broch. Admin. Fin.
AutothinkRAG 44.86 52.20 52.26 52.34 51.00 62.19 56.14 51.29
Nothink 43.15 48.55 43.21 52.25 46.00 53.08 50.42 47.21
VLM (direct) 41.43 49.27 43.43 52.25 41.00 54.32 48.72 45.55
VLM (unguided) 42.07 51.12 45.24 52.34 43.00 55.78 50.13 47.21
baseline 40.41 47.82 42.42 50.96 42.00 51.85 46.15 44.86
8B 42.80 44.12 43.21 52.34 46.00 50.61 50.42 46.09
RAG-Anything 41.09 41.30 44.22 45.16 40.00 54.32 52.13 44.36
@@ 3.3.2 Functional Decoupling via DPR
Recent studies indicate that VLMs exhibit significantly lower capabilities in logical reasoning compared to LLMs, particularly in document question answering tasks involving unanswerable queries. As reported in MMGraphRAG Wan and Yu (2025), LLM-based systems consistently outperform multimodal models on the DocBench and MMLongBench benchmarks. Notably, VLM-based models tend to produce hallucinated answers when the required evidence is absent (as shown in Table 1). In long-document multimodal QA, visual perception, cross-page evidence alignment, and reasoning are tightly coupled within a single generation, making the model prone to cascading errors: weakly aligned evidence causes overcommitment to unsupported answers instead of reliable abstention.Conversely, VLMs demonstrate superior performance in descriptive tasks post-perception. Consequently, we propose the Decoupled Perception and Reasoning (DPR) strategy, partitioning the task into two specialized sub-processes:
@@ 4 Experiments
In this section, we empirically validate the effectiveness of AutothinkRAG. Our goal is to address two primary research questions: (1) Compared to static retrieval, can the Autothink router effectively reduce hallucinations and handle complex query intents? (2) Does the Decoupled Perception-Reasoning (DPR) architecture demonstrate superior performance in reasoning tasks involving multimodal documents of varying lengths?
@@ 4.2 Implementation Details
In our experimental evaluation of AutothinkRAG, we implement a multi-model collaborative framework characterized by a decoupled architecture. Specifically, we employ Qwen2.5-VL-3B as the Autothink Router to facilitate global task orchestration and intent scheduling. For the knowledge construction phase, Qwen3-8B (non-reasoning version) is utilized to perform robust entity and relation extraction. To optimize retrieval precision, we adopt Qwen3-embedding-8B (with 1024-dimensional embeddings) for dense vector representation, followed by bge-reranker-v2-m3 for fine-grained reranking. To realize the Decoupled Perception-Reasoning (DPR) paradigm, we deploy Qwen2.5-VL-7B as the Perception Expert, which distills complex visual information into dense textual descriptions. Subsequently, Qwen3-32B-Instruct serves as the Reasoning Expert to execute high-level logical deductions based on the integrated context. Regarding hyperparameter configurations, all generative models are set with a temperature of $0.7$ and a Top-$p$ value of $0.9$. During the retrieval stage, we set the Top-$K$ for entity recall to $40$ and the number of retrieved text chunks to $20$ to ensure sufficient context coverage.
@@ 4.3 Main Results
Table 2 shows that AutothinkRAG achieves the best Overall Accuracy of 82.13%, outperforming the baseline (78.02%). The largest gain appears in Unanswerable (Una), where accuracy improves from 52.80% to 81.25% (+28.45%), indicating the effectiveness of the Autothink router in detecting insufficient evidence and reducing hallucinations. Consistent improvements on News (+10.83%) and Government (+8.30%) further demonstrate the advantage of Hybrid Hypergraph Retrieval in modeling complex entity relations beyond simple vector search.
@@
Table 3 reports the results on long-context tasks. AutothinkRAG consistently outperforms the baseline across all document categories, achieving an Overall Accuracy of 51.29% (+6.43%). While the baseline struggles with visual noise in long documents, our DPR Architecture remains effective in complex categories such as Admin (+10.34%) and Finance (+9.99%), bridging the “Reasoning Gap” by converting visual signals into textual descriptions and delegating synthesis to a specialized LLM.
To quantify the efficiency of AutothinkRAG, we analyze per-query token consumption and monetary cost (Table 4 and Table 5). The overall inference comprises three stages: query routing (Qwen2.5-VL-3B), visual description generation (Qwen2.5-VL-7B), and final answer synthesis (LLM). The input context of the synthesis stage accounts for the dominant cost. Without routing, pure hypergraph retrieval feeds substantial redundant context into the LLM, costing 63.0k tokens / 0.132 RMB per query. With the Autothink router, the synthesis input is significantly reduced to 51.4k tokens / 0.108 RMB, yielding savings of 18.9% and 18.2%, respectively. This demonstrates that the router effectively avoids unnecessary long-context reasoning by dynamically selecting retrieval paths based on query complexity.
@@ 4.4 Baselines
To ensure experimental rigor and fairness, we built a baseline model and employed the identical model on RAG-Anything (Guo et al., 2025) for testing and comparison.In the baseline configuration, we bypass query decomposition and employ the VLM for direct end-to-end inference, rather than using our proposed decoupled paradigm. For the RAG-Anything benchmark, we substitute its internal components with the specific models used in our framework and maintain identical hyperparameter configurations to ensure a controlled comparison.
@@ 4.5 Ablation Study
Pipeline Image Description Summary Total
Hypergraph 2.7k 60.3k 63.0k
Routing-Enhanced 2.7k 48.3k 51.4k
Savings (Abs.) - 12.0k 11.6k
Savings (Rel.) - 19.9% 18.9%
Table 4: Token comparison across stages.
Pipeline Image Description Summary Total
Hypergraph 0.0084 0.1239 0.132
Routing-Enhanced 0.0084 0.0984 0.108
Savings (Abs.) - 0.0255 0.024
Savings (Rel.) - 20.6% 18.2%
Table 5: Monetary cost comparison across stages.
@@
Our decoupling module transforms the direct reasoning of the VLM into a hybrid VLM+LLM indirect reasoning paradigm. To evaluate its efficacy, we conduct an ablation study by replacing this module with a standalone LLM reasoning component while keeping the upstream router constant. We evaluate the accuracy across various page ranges, as illustrated in Figure  4. The results indicate that both direct VLM reasoning and unguided VLM decoupling for visual perception in long-document settings substantially underperform the proposed full decoupling framework, with the performance gap becoming more pronounced as the document length increases.
@@ 5 Conclusion
While AutothinkRAG achieves significant performance gains over existing multimodal GraphRAG architectures using models with fewer parameters, it still relies on a sequential document parsing and embedding pipeline, which limits the overall processing speed. Our future work will be dedicated to achieving the efficient coupling of document parsing and information encoding to improve the overall processing speed and accuracy.